# Explanation as a Watermark: Towards Harmless and Multi-bit Model Ownership Verification via Watermarking Feature Attribution

### v1

@@ I Introduction
Ownership verification is currently the most critical and widely adopted method to safeguard model copyright. Specifically, it intends to identify whether a given suspicious third-party model is an unauthorized copy from model owners. Implanting owner-specified watermarks ($i.e.$, model watermarks) into the (victim) model is the primary solution for ownership verification [50]. Model watermarking methods generally have two main stages, including watermark embedding and ownership verification. In the first stage, model owners should embed a specific secret pattern ($i.e.$, watermark) that will be ‘inherited’ by unauthorized model copies into the model. After that, in cases where adversaries may illegally steal the victim model, the model owner can turn to a trusted authority for verification by examining whether the suspicious model has a similar watermark to the one implanted in the victim model. If this watermark is present, the suspect model is an unauthorized version of the victim model.
@@
The remaining parts of this paper are structured as follows: Section II provides an introduction to the background knowledge of DNN, Explainable Artificial Intelligence (XAI), and ownership verification of DNNs. Section III presents the problem formulation related to black-box DNN ownership verification. In Section IV, we propose a comprehensive insight, methodology, and implementation approach for EaaW. The effectiveness of EaaW is evaluated through experiments on image classification models and text generation models in Section V. Furthermore, additional experiments are conducted in Section VI to investigate the impact of EaaW on the watermarked models and analyze the security against ambiguity attacks. We conclude the paper in Section VII at the end.
@@ II-A Deep Neural Networks
While training a DNN model $\mathcal{M}$, the model takes the raw training data $\bm{x}\in\mathbb{R}^{m}$ as input, and then maps ${\bm{x}}$ to the output prediction ${\bm{p}}\in\mathbb{R}^{n}$ through a parametric function $\bm{p}=f(\bm{x};\Theta)$. The parametric functions $f(\cdot)$ are defined by both the architecture of the DNN model and the parameters $\Theta$. The developer then defines the loss function $\mathcal{L}(\cdot)$ to measure the difference between the output prediction $\bm{p}$ of the model and the true label $\bm{y}$. The objective of training the DNN model is equivalent to optimizing the parameters $\Theta$ to have the minimum loss. The optimization problem can be formally defined as Eq. (1).
@@
Recently, researchers have found that scaling the size of the models and datasets can make significant improvements in the models’ performance on various tasks, which leads to the rise of the large foundation models [52, 42]. Those large models often have more than 70 billion parameters and building such a large model requires a huge amount of training data, computational resources, and human expertise. Therefore, it is essential to protect the intellectual property rights of the model owners on the models.
@@ II-B Explainable Artificial Intelligence
There are three main categories of XAI techniques based on the application stages, $i.e.$, pre-modeling explainability, explainable modeling, and post-modeling explainability [39]. In this paper, we mainly focus on a specific type of post-modeling explainability method, feature attribution, in XAI [48, 46]. Feature attribution is a method that helps users understand the importance of each feature in a model’s decision-making process. It calculates a real-value importance score for each feature based on its impact on the model’s output. The score could range from a positive value that shows its contribution to the prediction of the model, a zero that would mean that the feature has no contribution, to a negative value that means that removing that feature would increase the probability of the predicted class.
@@ II-C Ownership Verification of DNN Models
Ownership verification of DNN models involves verifying whether the suspicious model is a copy of the model developed by another party (called the victim model[50]. Watermark [1, 5, 30] and fingerprint [4, 22, 56] are two different solutions to implementing ownership verification, while the watermark is the most widely used ownership verification technique. Model watermark refers to embedding a unique signature ($i.e.$, watermark), which represents the identity of the model owner, into the model [53, 1, 25]. The model owner can extract the watermark from the model in case the model is illegally used by the adversary.
In general, model watermarking methods can be divided into two categories, white-box model watermarking methods and black-box model watermarking methods, as follows.
White-box Model Watermarking Methods: white-box model watermarking methods embed the watermark directly in the model parameters [53, 35, 54]. For instance, Uchida et al. propose to add a watermark regularization term into the loss function and embed the watermark through fine-tuning [53]. White-box model watermarking methods assume that the verifier can have full access to the suspicious model during verification. This assumption is difficult to realize in practical scenarios because the model is usually black-box in the real world. Such a limitation prevents the application of the white-box watermarking methods. The watermark can also be embedded into the model via adjusting the architecture of the model [11, 35], or embedding external features [29, 28].
@@
However, because the backdoor-based model watermarking methods embed a zero-bit watermark into the prediction of the models, they incur several disadvantages. First, although backdoor-based methods claim that they do not significantly compromise the functionality of the model with benign datasets, backdoor-based model ownership verification can still be harmful. Second, backdoor-based methods are based on misclassification and can only identify the presence or absence of a watermark. Adversaries can easily manipulate adversarial samples to verify their ownership on the victim model, leading to ambiguity in ownership verification [32]. To the best of our knowledge, BlackMarks [3] is the only multi-bit black-box watermarking method, based on the harmful backdoor attack. BlackMarks divides the output classes of the model into two groups. If the prediction class of the $i-$th trigger sample belongs to the first group, it means the $i-$th bit in the watermark is 00, otherwise $1$. BlackMarks makes the sequential predictions of the trigger samples as a multi-bit watermark. However, the adversaries can create any bit string by rearranging the input trigger samples, leading to ambiguity.
Moreover, Maini et al. proposed a non-backdoor black-box model watermarking method called Dataset Inference (DI) [36, 10]. However, some recent studies demonstrated that DI may make misjudgments [29]. DI is also not able to embed a multi-bit watermark into the model. These limitations hinder its applicability in practice. How to design an effective and multi-bit black-box model watermarking method remains an important open question.
@@ III-A Threat Model
  • The adversary can conduct several watermark removal attack techniques trying to remove the watermark in the victim model, such as the fine-tuning attack, the model pruning attack, and the adaptive attack.
@@ III-B Design Objectives
  • Effectiveness: Effectiveness signifies that the watermark needs to be properly embedded into the model. If the suspicious model actually originates from the victim model, the ownership verification algorithm can deterministically output a watermark which is similar to the victim’s pre-designed watermark.
@@ IV-A Insight and Overview of EaaW
As discussed in Section II-C, backdoor-based model watermarks encounter two-fold drawbacks and challenges, namely harmfulness and ambiguity. These drawbacks arise primarily from the ‘zero-bit’ nature of the backdoor-based methods, where the watermark is embedded into the binary status of the model predictions. To tackle these challenges, a crucial question arises: ‘Can we find an alternative space to embed a multi-bit watermark without changing the predictions?’ Inspired by eXplainable Artificial Intelligence (XAI) and feature attribution, we propose EaaW. Our primary insight is that instead of directly watermarking the prediction classes of the model, the explanation generated by feature attribution algorithms can also serve as a suitable carrier for hiding information and embedding watermarks.
@@ IV-C Watermark Extraction through Feature Attribution
After generating the masks, we construct the masked samples $\mathcal{X}_{m}$ to constitute a dataset by masking the basic parts in the trigger sample according to the randomly generated masks. The masking operation can be denoted as $\otimes$.
$\mathcal{X}_{m}=M\otimes\mathcal{X}_{T}.$ (4)
Specifically, if the element in the mask $M_{i}$ is $1$, the corresponding basic part preserves the original value. Otherwise, the corresponding basic part should be replaced by a certain value if the element is 00. For example, for the image data, the pixels in the masked basic part could be aligned by 00. The tokens in the text data can be replaced by a special token ‘<unk>’which represents an unknown token. The examples of the masked samples are shown in Figure 2.
@@
The metric function $\mathcal{M}(\cdot)$ needs to be derivable and can provide a quantificational evaluation of the output. Users can customize it based on the specific DL task and prediction form. Since there usually exists a derivable metric function in DL tasks ($e.g.$, loss function), EaaW can easily be extended to various DL tasks such as image classification and text generation. The implementation details and results are shown in Section V-A and Section V-B, respectively.
@@ V Experiments
V-A Results on Image Classification Models
V-A1 Experimental Settings
@@ V-A2 Evaluation on Effectiveness and Harmlessness
Figure 3 and Table I present the experimental results of watermarking image classification models, demonstrating the successful embedding of the multi-bit watermark into these models via the utilization of EaaW. The p-values are far less than the significant level $\alpha$ and the WSRs are nearly equal to $1$. Those results unequivocally establish the effectiveness of EaaW in facilitating watermark embedding.
@@ V-A3 Evaluation on Distinctiveness
To evaluate the distinctiveness of EaaW, we also carry out experiments to test whether the watermark can be extracted with independently trained models and independently selected trigger samples. The results are shown in Table II, indicating that the watermark extracted with independent models and independent triggers cannot pass the ownership verification process. The minimal p-value with independent models and independent triggers is $0.012$ which is still $>0.01$ and far greater than the p-value of the watermarked model. The WSRs with independent models and triggers are mostly near $50\%$. Furthermore, the results also indicate that incorporating a higher number of bits in the watermark enhances distinctiveness and security. This is evidenced by the smaller p-values obtained when extracting a fewer-bit watermark using an independently trained model.
An example of the extracted watermarks with the watermarked model, independent model, and the independent trigger is shown in Figure 4. It shows that the extracted watermarks in the latter two scenarios are completely different from the original watermark, further proving the distinctiveness.
@@ V-B Results on Text Generation Models
Figure 4: An example of the extracted watermarks with the watermarked model, independent model, and independent trigger.
@@ V-B1 Experimental Settings
In this section, we adopt EaaW to watermark the text generation models. The text generation model, especially the casual language model, is a type of language model that predicts the next token in a sequence of tokens [44]. Text generation models are widely used as the pre-trained foundation models in various natural language processing (NLP) tasks [42].
We take GPT-2 [44] as an example to evaluate EaaW on the text generation model since it is a representative open-sourced transformer-based model and many state-of-the-art large language models have similar structures. The experiments with another popular text generation model, $i.e.$, BERT [7], can be found in Appendix -C. Four different datasets, including wikitext [38], bookcorpus [61], ptb-text-only [37], and lambada [43], are used to fine-tune the GPT-2 model and embed the multi-bit watermark. We randomly select a sequence in the training set as the trigger sample. We also randomly generate a $k$-bit string as the watermark and the examples of the text trigger samples and the embedded watermarks can be found in Appendix -A. The lengths of the watermark are set to $32$, $48$, $64$, $96$, and $128$. Additionally, different from the image classification model, we utilize the average prediction probabilities of the target tokens in the masked trigger sequence as the metric vector $\bm{v}$. The implementation details can be found in Appendix -A.
@@ V-B3 Evaluation on Distinctiveness
Similar to the experiments conducted on the image classification models, we also test whether the watermark can be extracted from the independently trained language model or with independent trigger samples to validate the distinctiveness of EaaW.
From Table IV, we can find that the p-values with the independent model or independent trigger are greater than the significant level $\alpha=0.01$, and the WSRs are around $0.5$, which is similar to the results of the experiments on the image classification model. In addition, we can find that when the length of the watermark is relatively small, e.g. $32$, the p-value with the independent model is small with a minimum $0.013$ and close to the significant level. As the length of the watermark increases, the p-value with the independent model also increases, suggesting that embedding a watermark with more bits can obtain better security and distinctiveness.
@@ V-C1 The Resistance to Fine-tuning Attack
Fine-tuning refers to training the watermarked model with a local benign dataset for a few epochs. In the fine-tuning attack, the adversary may attempt to remove the watermark inside the model through fine-tuning. We fine-tune the EaaW-watermarked models with 20 epochs wheretraining has converged.
Figure 5 shows the log p-value and the WSR during the fine-tuning attack, while Figure 12 shows the extracted watermarks before and after the fine-tuning attack. As shown in Figure 5, the p-value and WSR fluctuate during fine-tuning, whereas the p-value is always significantly lower than the significant level $\alpha$ (denoted by the purple dotted line) and the WSR is greater than $0.85$. These results demonstrate the resistance of our EaaW to fine-tuning attacks. We argue that this is mostly because we did not change the label of watermarked samples during model training, leading to minor effects of fine-tuningcompared to backdoor-based methods.

### v2

@@ Abstract
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2025 23 - 28 February 2025, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2025.23338 www.ndss-symposium.org
@@ I Introduction
Ownership verification is currently the most critical and widely adopted method to safeguard model copyright [55, 37]. Specifically, it intends to identify whether a given suspicious third-party model is an unauthorized copy from model owners. Implanting owner-specified watermarks ($i.e.$, model watermarks) into the (victim) model is the primary solution for ownership verification [55]. Model watermarking methods generally have two main stages, including watermark embedding and ownership verification. In the first stage, model owners should embed a specific secret pattern ($i.e.$, watermark) that will be ‘inherited’ by unauthorized model copies into the model. After that, in cases where adversaries may illegally steal the victim model, the model owner can turn to a trusted authority for verification by examining whether the suspicious model has a similar watermark to the one implanted in the victim model. If this watermark is present, the suspect model is an unauthorized version of the victim model.
@@ II-A Deep Neural Networks
While training a DNN model $\mathcal{M}$, the model takes the raw training data $\bm{x}\in\mathbb{R}^{m}$ as input, and then maps ${\bm{x}}$ to the output prediction ${\bm{p}}\in\mathbb{R}^{n}$ through a parametric function $\bm{p}=f(\bm{x};\Theta)$. The parametric functions $f(\cdot)$ are defined by both the architecture of the DNN model and the parameters $\Theta$. The developer then defines the loss function $\mathcal{L}(\cdot)$ to measure the difference between the output prediction $\bm{p}$ of the model and the true label $\bm{y}$. The objective of training the DNN model is equivalent to optimizing the parameters $\Theta$ to have the minimum loss, which can be formally defined as Eq. (1).
@@ II-B Explainable Artificial Intelligence
There are three main categories of XAI techniques based on the application stages, $i.e.$, pre-modeling explainability, explainable modeling, and post-modeling explainability [42]. In this paper, we mainly focus on a specific type of post-modeling explainability method, feature attribution, in XAI [53, 51]. Feature attribution is a method that helps users understand the importance of each feature in a model’s decision-making process. It calculates a real-value importance score for each feature based on its impact on the model’s output. The score could range from a positive value that shows its contribution to the prediction of the model, a zero that means the feature has no contribution, to a negative value that implies removing that feature could increase the probability of the predicted class.
@@ II-C Ownership Verification of DNN Models
Ownership verification of DNN models involves verifying whether the suspicious model is a copy of the model developed by another party (called the victim model[55]. Watermark [1, 6, 33] and fingerprint [5, 23, 62] are two different solutions to implementing ownership verification. Model watermark refers to embedding a unique signature ($i.e.$, watermark), which represents the identity of the model owner, into the model [1, 28]. The model owner can extract the watermark from the model in case the model is illegally used by the adversary. In general, model watermarking methods can be divided into two categories, white-box and black-box model watermarking methods, as follows.
White-box Model Watermarking Methods: white-box model watermarking methods embed the watermark directly in the model parameters [58, 38, 60]. For instance, Uchida et al. proposed to add a watermark regularization term into the loss function and embeded the watermark through fine-tuning [58]. The watermark can also be embedded into the model via adjusting the architecture of the model [12, 38], embedding external features [32, 31], or introducing a transposed model [26]. White-box model watermarking methods assume that the verifier can have full access to the suspicious model during verification. This assumption is difficult to realize in practical scenarios because the model is usually black-box in the real world. Such a limitation prevents the application of the white-box watermarking methods.
@@
However, because the backdoor-based model watermarking methods embed a zero-bit watermark into the prediction of the models, they incur several disadvantages. First, although backdoor-based methods claim that they do not significantly compromise the functionality of the model with benign datasets, backdoor-based model ownership verification can still be harmful. Second, backdoor-based methods are based on misclassification and can only identify the presence or absence of a watermark. Adversaries can easily manipulate adversarial samples to verify their ownership on the victim model, leading to ambiguity in ownership verification [35].
To the best of our knowledge, BlackMarks [4] is the only multi-bit black-box watermarking method, based on the harmful backdoor attack. BlackMarks divides the output classes of the model into two groups. If the prediction class of the $i-$th trigger sample belongs to the first group, it means the $i-$th bit in the watermark is 00, otherwise $1$. BlackMarks makes the sequential predictions of the trigger samples as a multi-bit watermark. However, the adversaries can create any bit string by rearranging the input trigger samples, leading to ambiguity. Moreover, Maini et al. proposed a non-backdoor black-box model watermarking method called Dataset Inference (DI) [39, 11]. However, some recent studies demonstrated that DI may make misjudgments [32]. DI is also not able to embed a multi-bit watermark into the model. These limitations hinder its applicability in practice.
Model Fingerprinting Methods: Model fingerprinting methods provide another solution for model ownership verification. Model fingerprinting aims to identify the intrinsic feature ($i.e.$, fingerprint) of the model. By comparing the fingerprints of two models, we can judge whether one model is a copy of the other. In general, model fingerprinting methods can be categorized into two types. The first is based on adversarial examples (AE) [3, 48]. AE-based methods exploit adversarial examples to characterize the decision boundary of the model. The other type is the testing-based methods [5, 23], which compares the outputs of the two models on a specific mapping function. Although model fingerprinting methods do not need to modify the model, they are not always effective in distinguishing models, especially under attacks [46]. Besides, model fingerprinting methods cannot embed any identity information and are also vulnerable to ambiguity attacks [35].
II-D Watermark Removal Attack
The adversaries may adopt watermark removal attacks to remove the watermark of victim models to evade ownership verification. Generally, existing watermark removal attacks can be categorized into two types: unintentional removal attacks and intentional adaptive attacks [37].
On the one hand, some model reuse techniques may unintentionally remove the watermark in the model. These techniques include fine-tuning and model pruning [17]. On the other hand, if the adversary knows the watermarking method, it can adaptively design the removal attacks. There are two representative adaptive attacks, namely the overwriting attack and the unlearning attack [37]. The former tries to embed another watermark into the model to overwrite the original one, while the latter aims to unlearn the watermark by updating the model in the direction opposite to the watermark gradient.
@@ III-A Threat Model
  • The adversary can conduct several watermark removal attack techniques trying to remove the watermark in the victim model, such as the fine-tuning attack and the model pruning attack. The adversary may also be aware of the watermarking technique and can carry out adaptive attacks to remove the watermark.
@@ III-B Design Objectives
  • Effectiveness: Effectiveness signifies that the watermark needs to be properly embedded into the model. If the suspicious model actually originates from the victim model, the ownership verification algorithm can deterministically output a watermark that is similar to the victim’s pre-designed watermark.
@@ IV-A Insight and Overview of EaaW
As discussed in Section II-C, backdoor-based model watermarks encounter two-fold drawbacks and challenges, namely harmfulness and ambiguity. These drawbacks arise primarily from the ‘zero-bit’ nature of the backdoor-based methods, where the watermark is embedded into the binary status of the model predictions. To tackle these challenges, a crucial question arises: ‘Can we find an alternative space to embed a multi-bit watermark without changing the predictions?’ Inspired by XAI and feature attribution, we propose EaaW. Our primary insight is that instead of directly watermarking the prediction classes of the model, the explanation generated by feature attribution algorithms can also serve as a suitable carrier for hiding information and embedding watermarks.
@@ IV-C Watermark Extraction through Feature Attribution
After that, we construct the masked samples $\mathcal{X}_{m}$ to constitute a dataset by masking the basic parts in the trigger sample according to the randomly generated masks. The masking operation can be denoted as $\otimes$, $i.e.$, $\mathcal{X}_{m}=M\otimes\mathcal{X}_{T}$. Specifically, if the element in the mask $M_{i}$ is $1$, the corresponding basic part preserves its original value. Otherwise, the basic part is replaced by a certain value if the element is 00. The examples of the masked samples are shown in Figure 2.
@@
The metric function $\mathcal{M}(\cdot)$ needs to be derivable and can provide a quantificational evaluation of the output. Users can customize it based on the specific DL task and prediction form. Since there usually exists a derivable metric function in DL tasks ($e.g.$, loss function), EaaW can easily be extended to various DL tasks.
@@
Implementation Examples. From the above introduction, the key to applying the watermark extraction to different tasks is to design the rule of the masking operation $\otimes$ and the metric function $\mathcal{M}(\cdot)$. The masking operation is used to construct the masked samples and the metric function measures the quality of the predictions. We hereby present two implementation examples of image classification models and text generation models, as shown in Figure 2. Specifically, for image classification models, the masking operation aligns the pixels in the masked basic part by 00 and keeps the original values of other pixels. Additionally, we can choose the function outputting the predicted probability of the ground-truth class as the metric function; For text generation models, especially the casual language model [45], the masking operation replaces the masked tokens with a special token ‘<unk>’, which represents an unknown token. We exploit the function generating the average prediction probabilities of the target tokens in the masked trigger sequence as the metric function.
@@ V Experiments
V-A Results on Image Classification Models
V-A1 Experimental Settings
@@ V-A2 Evaluation on Effectiveness and Harmlessness
Figure 3 and Table I present the experimental results of watermarking image classification models, demonstrating the successful embedding of the multi-bit watermark into these models via the utilization of EaaW. The p-values are far less than the significant level $\alpha$ and the WSRs are nearly equal to $1$. Those results unequivocally establish the effectiveness of EaaW in facilitating watermark embedding. Besides, since the WSR is nearly $1$, the statistic in the chi-squared test is approximately proportional to $1/k$, where $k$ is the length of the watermark [50]. As such, the p-value decreases when $k$ increases.
@@ V-A3 Evaluation on Distinctiveness
To evaluate the distinctiveness of EaaW, we also carry out experiments to test whether the watermark can be extracted with independently trained models and independently selected trigger samples. We exploit the independently trained ResNet-18 that does not have a watermark as the independent model and we use the other trigger samples as the independent trigger samples. The results are shown in Table II, indicating that the watermark extracted with independent models and independent triggers cannot pass the ownership verification process. The minimal p-value with independent models and independent triggers is $0.012$ which is still $>0.01$ and far greater than the p-value of the watermarked model. The WSRs with independent models and triggers are mostly near $50\%$. Furthermore, the results also indicate that incorporating a higher number of bits in the watermark enhances distinctiveness and security. This is evidenced by the smaller p-values obtained when extracting a fewer-bit watermark using an independently trained model.
@@ V-B1 Experimental Settings
In this section, we adopt EaaW to watermark the text generation models. The text generation model, especially the casual language model, is a type of language model that predicts the next token in a sequence of tokens [49]. Text generation models are widely used as the pre-trained foundation models in various tasks [45].
We take GPT-2 [49] as an example to evaluate EaaW on the text generation model since it is a representative open-sourced transformer-based model and many state-of-the-art large language models have similar structures. The experiments with another popular text generation model, $i.e.$, BERT [8], can be found in Appendix -C. Four different datasets, including wikitext [41], bookcorpus [67], ptb-text-only [40], and lambada [47], are used to fine-tune the GPT-2 model and embed the multi-bit watermark. We randomly select a sequence in the training set as the trigger sample. We also randomly generate a $k$-bit string as the watermark and the examples of the text trigger samples and the embedded watermarks can be found in Appendix -B. The lengths of the watermark are set to $32$, $48$, $64$, $96$, and $128$. Additionally, different from the image classification model, we utilize the average prediction probabilities of the target tokens in the masked trigger sequence as the metric vector $\bm{v}$. The implementation details can be found in Appendix -B.
@@ V-B3 Evaluation on Distinctiveness
Similar to the experiments conducted on the image classification models, we also test whether the watermark can be extracted from the independently trained language model or with independent trigger samples to validate the distinctiveness of EaaW. We also exploit the model without the watermark as the independent model and randomly choose several sequences of tokens as the independent trigger samples.
From Table IV, we can find that the p-values with the independent model or independent trigger are greater than the significant level $\alpha=0.01$, and the WSRs are around $0.5$, which is similar to the results of the experiments on the image classification model. In addition, we can find that when the length of the watermark is relatively small, e.g. $32$, the p-value with the independent model is small with a minimum of $0.013$ and close to the significant level. As the length of the watermark increases, the p-value with the independent model also increases, suggesting that embedding a watermark with more bits can obtain better security and distinctiveness.
@@ V-C1 The Resistance to Fine-tuning Attack
Fine-tuning refers to training the watermarked model with a local benign dataset for a few epochs. In the fine-tuning attack, the adversary may attempt to remove the watermark inside the model through fine-tuning. We fine-tune the EaaW-watermarked models with 20 epochs on the testing set where the training has converged.
Figure 4 shows the log p-value and the WSR during the fine-tuning attack. The p-value and WSR fluctuate during fine-tuning, whereas the p-value is always significantly lower than the significant level $\alpha$ (denoted by the purple dotted line) and the WSR is greater than $0.85$. These results demonstrate the resistance of our EaaW to fine-tuning attacks. We argue that this is mostly because we did not change the label of watermarked samples during model training, leading to minor effects of fine-tuning compared to backdoor-based methods.